What will be the volume of 0.01 M HCl is required to completely react with 4g NaOh

Grade:12th Pass

1 Answers

Arun
5 years ago

Dissociation of the two reactants is as below:

NaOH -> Na+ + OH-

HCl -> H+ + Cl-

( Note: 1 OH- neutralizez 1 H+ ion)

Now, No. of moles of NaOH = Given mass of substance/ molar mass = 4 g / 23g+16g = 4/40 = 0.1 mol

NaOH will give 0.1 mol of OH- ions.

Therefore, o.1 mol of H+ ions will be required to neutralize it.

Now using molarity formula,

Molarity of HCl solution = No. of moles of HCl/ Vol. of solution in Litres

0.01 = 0.1 / V

V = 0.1/0.01 = 10 L = 10 Litres

So, the volume of 0.35 M HCl solution is 10 Litres.